Question 13 Determine the domain of f(x)= x2 - 4x +3 x-1 in interval notation.
Question 14 Given that f(x) = - 4x + 3 is a one-to-one function. Determine f-(x).
Answer #1

13 f(x) = x² - 4x +3 x-1 f(x) = (x-3) (x-1) 2 x X-3 f(x) Interval notation .(-00 ,, 1) u (1,0) 14 f(x) = - 4x + 3 y = - 4x +3Plotting: y 2 x - 4x + 3 X-1 5 -5 5 10 -5.